# Intrinsic exceptional point – a challenge in quantum theory

### v1

@@ 5 Towards a regularization of $H^{(IEP)}$s by perturbation
From a purely historical point of view the idea of “prominence” of the IEP-sampling Schrödinger Eq. (1) dates back to its methodical role in field theory [21] and to the Bessis’ and Zinn-Justin’s empirically revealed conjecture (cf. [4], cited also in [5]) that the spectrum $\{E_{n}^{(IC)}\}$ of $H^{(IC)}$ is real, discrete and bounded from below, i.e., tractable, in principle at least, as the set of the observable energy levels of bound states of a unitary quantum system. In spite of the manifest non-Hermiticity of Hamiltonian $H^{(IC)}$, the model was temporarily accepted as potentially compatible with all of the principles and postulates of quantum mechanics. The corresponding technical details may be found in review paper [22].
Unfortunately, the end of the excitement came after the Siegl’s and Krejčiřík’s rigorous proof that the IC model cannot in fact be assigned any form of conventional probabilistic interpretation in a mathematically consistent manner [1]. More or less the same conclusion has been also made, in a not yet published preprint, by Günther and Stefani [8]. At present, in the context of the unitary-evolution part of non-Hermitian quantum mechanics the problem of a correct physical interpretation of the IC model (1) remains, therefore, unresolved.
@@ 5.1 The IEP - EPN differences and parallels
The essence of the anomalous nature of any IEP-related Hamiltonian $H^{(IEP)}$ lies in the asymptotic degeneracies (23) and (24) of its respective right and left eigenvectors. At the same time, a weak point of the amendment of the basis as mediated by the choice of non-diagonal matrix (26) may be seen in the necessity of specification of a “sufficiently large” onset $K\gg 1$ of the de-parallelization. Such a specification is just numerically, computer-precision motivated. In contrast to the above-outlined treatment of the EPN scenarios where the dimension $N$ was fixed, the IEP-implied choice of any finite $K$ is purely pragmatic and virtually arbitrary.
We now intend to show that, surprisingly enough, the apparently more or less accidental flexibility of our choice of $K$ can in fact become an important tool facilitating an EPN-resembling regularization and consistent interpretation of quantum systems near their IEP singular extreme.
First of all, there is no doubt about the necessity of transition from the less suitable unperturbed basis (composed of eigenvectors) to an “anomalous” basis resembling Eq. (4). The reason is the situation as represented by Eqs. (25) and (26): only a rectification of the underlying biorthogonal or biorthonormal basis [23] can re-establish the EPN - IEP parallels even when achieved, also in the latter case, at an expense of the non-diagonality and non-Hermiticity of matrix (26).
Although the EPN singularity encountered at finite matrix dimensions $N<\infty$ is, according to paper [1], perceivably weaker than its IEP analogue, the essence of our present message will be complementary. Basically, we will emphasize that one can also reveal and make a productive use of certain partial similarities between the two concepts. In particular, we propose that the above-outlined possibility and feasibility of treating the manifestly unphysical finite-dimensional singular matrices $H^{(N)}(g^{(EPN)})$ as formally acceptable unperturbed Hamiltonians is to be transferred also to the IEP context. Moreover, also the construction of the equally anomalous EPN-related “good” basis (4) composed of the columns of the transition matrix should be, mutatis mutandis, introduced and reconstructed from any given anomalous Hamiltonian $H^{(IEP)}$.
@@ 5.2 IEP-unfolding bases
${\cal R}^{(IEP)}=\{|\psi_{0}\rangle,|\psi_{1}\rangle,\ldots,|\psi_{K-1}\rangle,|f_{K}\rangle,|f_{K+1}\rangle,|f_{K+2}\rangle,\ldots\}$ (27)
composed of the mere first $K$ eigenkets $|\psi_{j}\rangle$ complemented by the modified, associated-like ket vectors $|f_{K+k}\rangle$ with $k=0,1,2,\ldots$.
@@ 5.3 Recurrences
Our freedom of choice of the highest-component coefficients $c_{k,k}$ enables us to suppress the IEP-accompanying asymptotic parallelization of the vectors of basis in Hilbert space. The goal is achieved. For every particular IEP model we recall recurrences (30) and replace the $Q-$projected part of the basis composed of eigenvectors by the $Q-$projected part of the basis composed, up to the first item $|f_{K}\rangle=c_{0,0}|\psi_{K}\rangle$, of non-eigenvectors. And this is precisely what has been done in Eq. (27).
@@ 6.1 Formulation of the problem
Günther with Stefani [8] stated that “what is still lacking” in the IEP setup “is a simple physical explanation scheme for the non-Rieszian behavior of the eigenfunction sets”. We agree. We are even more skeptical because we would rather say that the expected ‘simple physical explanation’ making, in particular, the popular IC oscillators (1) physical need not exist at all. Indeed, we believe that a consistent physical closed quantum system interpretation could much more easily be assigned to suitable perturbations of the “seed” IEP oscillators with uncertain interpretation (cf. [24] in this respect).
@@
Incidentally, a similar simplification has also been detected in the realistic EPN-supporting Bose-Hubbard model of paper [12] where, in dependence on parameters, the authors had to use both  the degenerate and non-degenerate versions of perturbation theory. Thus, no abstract conceptual problems have to be expected to emerge after one returns to the generic IEP-related dynamics. Still, as long as the IEP-related problems are infinite-dimensional, the perturbed IEP spectrum cannot be deduced from any analogue of the EPN-based implicit-definition constraint (18). The study of properties of the vicinity of the IEP singularity cannot be based on a direct reference to its EPN analogue.
@@ 6.2 Structure of solutions
In the light of the $P+Q$ partitioning of matrix ${\cal J}^{(IEP)}$ of Eq. (26) the constructiton of the perturbed forms of the low-lying bound states remains standard. The $P-$projected states may be ignored just as certain decoupled observers. Only the treatment of the “asymptotic”, $Q-$projected components of the quantum system in question becomes difficult and singular, “for instance, due to spectral instabilities” [1].
@@
In our present, conceptually more ambitious analysis of the problem it makes sense to turn attention to the states with the high-lying zero-order unperturbed energies. The related unperturbed ket vectors $|\psi(0)\rangle$ will lie in the complementary (and infinite-dimensional) $Q-$projected subspace of Hilbert space. The more or less conventional construction of its perturbed descendant of Eq. (33) will then possess several anomalous features of course.
@@
Due to the specific upper-triangular two-diagonal structure of matrix $Q{\cal J}^{(IEP)}Q$, the construction of a biorthonormalized basis would be needed. Thus, the left eigenvectors of $Q{\cal J}^{(IEP)}Q$ (i.e., in the notation of paper [20], ketkets, $|\chi_{j}\rangle\!\rangle$) will be complicated and different from their right-eigenvector biorthogonal partners $|\chi_{j}\rangle$. This means that also the conventional Rayleigh-Schrödinger elementary unperturbed projectors $|\chi_{j}\rangle\!\rangle\langle\!\langle\chi_{j}|$ (needed during the construction) will have a practically prohibitively complicated explicit matrix structure.
@@
Another, independent word of warning might originate from the fact that for all of the truly high energy levels $E_{K+m}(\lambda)$ with $m\gg 1$ the use of the explicit Rayleigh-Schrödinger recipe would require the derivation of formulae which would be $m-$dependent and different for different $m-$numbered excitations. Fortunately, the latter technical obstacle and difficulty has a comparatively elementary resolution because what is fully at our disposal is our choice of the value of $K$. We may feel free to work, exclusively, with the properly innovated Rayleigh-Schrödinger formulae deduced just at a single value of $m$, i.e., say, at $m=0$.
@@
Similarly, from the second row of Eq. (39) we would then obtain, more naturally, the value of the second wave-function component
@@ 7 Discussion
All in all we can conclude that in both the EPN- and IEP-related unitary-evolution scenarios the properly amended form of perturbation theory seems to be able to provide, even in its leading-order form, some explicit and useful criteria of the acceptability or unacceptability of various preselected perturbations of potential phenomenological interest.
@@ 7.1 Benign perturbations
Between the EPN and IEP alternatives there still exists a crucial difference. Indeed, in the typical EPN-related analysis our considerations usually start from our knowledge of the “physical” family of models $H(g)$. Then, the only truly difficult problem is to localize the EPN singularity, especially when the values of $N$ are not too small. In the case of the IEP singularities, in contrast, we, typically, know only the unperturbed Hamiltonian as sampled here by the IC operator. It is possible to conclude that precisely this makes the IEP-related models perceivably more difficult to study.
From a purely pragmatic point of view a source of certain optimism could really be drawn even from the above-outlined leading-order perturbation-approximation criteria. Their strength certainly lies in the possibility of identification of the “malign” IEP perturbations which would destroy the reality of the spectrum and which would make the evolution non-unitary. The complementary problem of a reliable identification of the opposite, “benign” perturbations has to be declared to stay a mathematically much more difficult open question.
Incidentally, qualitatively the same conclusions have already been obtained in the simpler EPN context. For example, in the above-mentioned study [12] of a specific Bose-Hubbard model near its EPN dynamical extreme, the authors did not insist on the reality of the spectrum and treated their mathematical results as applicable in a broader, not necessarily unitary open-system context.
In a narrower, closed-system setting, a deeper analysis and resolution of the apparent EPN-related instability paradox was described in paper [18] in which we studied the exact as well as approximate secular equations in more detail. Our ultimate conclusion was that the necessary smallness condition specifying the class of the admissible, unitarity non-violating perturbations does not involve their upper-triangular matrix part at all. In contrast, for the perturbed-EPN model in question, the lower-triangular matrix part of all of the “benign” (i.e., unitarity-compatible) perturbations has been shown to have, obligatorily, the following matrix-element-dependent form of the condition of sufficient smallness in the $\lambda\to 0$ limit,
@@ 7.2 IC oscillator as popular toy model
In order to elucidate the benchmark-model role of the IC IEP oscillator from a historical perspective let us recall paper [5] in which Bender with Boettcher examined a rather broad family of time-independent non-Hermitian toy-model Hamiltonians (cf. Eq. (45) in Appendix A. 2 below). They felt guided by the postulate of (antilinear) symmetry of their models,
${\cal PT}H^{(BB)}=H^{(BB)}{\cal PT}\,$ (44)
where the linear operator ${\cal P}$ can be treated as parity (causing the space reflection $x\to-x$) while ${\cal T}$ has to mimic the antilinear time reversal.
These authors proposed to treat their operators $H^{(BB)}$ as “Hamiltonians whose spectra are real and positive” so that “these ${\cal PT}-$symmetric theories may be viewed as analytic continuations of conventional theories from real to complex phase space” [5]. During the subsequent wave of development of the related mathematics it has been revealed that in the language of functional analysis the ${\cal PT}-$symmetry (44) can be re-read as the so called pseudo-Hermiticity [14] as well as a self-adjointness in the Krein space with indefinite pseudo-metric ${\cal P}$ [25, 26, 27].
@@
This was also the feature which attracted a lot of attention among physicists [5, 22, 26, 28] precisely because they happened to generate the purely real, discrete and non-negative (i.e., hypothetically, observable and bound-state-like) energy spectra. Still, the ultimate verdict by mathematicians [1, 8] was discouraging because they proved that the IC Hamiltonian cannot be assigned any isospectral self-adjoint avatar $\mathfrak{h}({{t}})$ or acceptable physical inner-product metric [1]. Thus, the rigorous mathematical analysis finally led to the loss of some of the most optimistic phenomenological expectations.
@@ 8 Summary
Not too surprisingly, the highly desirable proofs of the so called unbroken form of ${\cal PT}-$symmetry (in which, by definition [22], the spectrum remains real) appeared to be, in numerous applications, strongly model-dependent. There seemed to be no universal criteria guaranteeing the existence of the unbroken ${\cal PT}-$symmetry in dependence on a suitable measure or degree of the non-Hermiticity of the Hamiltonian.
@@
In our present paper we explained that and how such a decision enabled us not only to pick up a rather natural measure of the non-Hermiticity (characterized simply by the inverse distance of $g$ from its exceptional value) but also to formulate a well-defined project in which we developed and applied, consequently, some innovative and suitable perturbation-approximation techniques. In its framework we managed to show that the unbroken ${\cal PT}-$symmetry of our models can really survive inside an open parametric domain, on a point of boundary of which our measure of non-Hermiticity reaches its maximum. The latter point has been shown to coincide either with the Kato’s [7] exceptional point of a finite order $N$ or with its hypothetical IEP analogue.
Such an approach has been found productive. Using certain slightly modified techniques of perturbation theory of linear operators we found that, paradoxically, the restriction of attention to the smallest vicinity of the singularity (in which the Hamiltonians become maximally non-Hermitian) leads to a remarkable simplification of the perturbation-approximation constructions. The special, “exceptional” non-diagonalizable operators appeared to be eligible as unperturbed Hamiltonians in spite of their being singular and unacceptable as observables. In their vicinity (denoted as ${\cal D}$) their diagonalizability as well as the observability status were re-established.
Constructively we managed to clarify also some of the consequences. The emergence of qualitative as well as quantitative EPN - IEP parallels helped us to complement and understand better the twelve years old disproof of the internal mathematical consistency of the IC IEP quantum oscillator [1]. Such a clarification can be perceived as being of a fundamental importance in quantum theory. Indeed, potentially, most of our observations might immediately be extended also to many other currently popular non-Hermitian quantum models.
@@
Fortunately, a temporary resolution of the conflict has been found in a rediscovery and return to a half-forgotten but still fully relevant older review paper by Scholtz et al [10]. In it, most of the objections by mathematicians were circumvented by an ad hoc  technical assumption that one is only allowed to consider the non-Hermitian Hamiltonians (as well as any other candidates for observables) which are, as operators in Hilbert space, bounded: cf. also several mathematically oriented reviews in [26] in this respect. Thus, one may call such a mathematically consistent version of the theory quasi-Hermitian quantum mechanics.

### v2

@@ 5 Towards a regularization of $H^{(IEP)}$s by perturbation
From a purely historical point of view the idea of “prominence” of the IEP-sampling Schrödinger Eq. (1) dates back to its methodical role in field theory [21] and to the Bessis’ and Zinn-Justin’s empirically revealed conjecture (cf. [4], cited also in [5]) that the spectrum $\{E_{n}^{(IC)}\}$ of $H^{(IC)}$ is real, discrete and bounded from below, i.e., tractable, in principle at least, as a set of observable energy levels. In spite of the manifest non-Hermiticity of Hamiltonian $H^{(IC)}$, the model was temporarily accepted as potentially compatible with all of the principles and postulates of quantum mechanics. The corresponding technical details may be found in review paper [22].
Unfortunately, the end of the excitement came after the Siegl’s and Krejčiřík’s rigorous proof that the IC model cannot in fact be assigned any form of conventional probabilistic interpretation in a mathematically consistent manner [1]. More or less the same conclusion has been also made, by Günther and Stefani, in a not yet published preprint [8]. At present, in the context of the unitary-evolution part of non-Hermitian quantum mechanics the problem of a correct physical interpretation of the IC model itself remains unresolved.
@@ 5.1 The IEP - EPN differences and parallels
The essence of the anomalous nature of any IEP-related Hamiltonian $H^{(IEP)}$ lies in the asymptotic degeneracies (23) and (24) of its respective right and left eigenvectors. At the same time, a weak point of the amendment of the basis as mediated by the choice of non-diagonal matrix (26) may be seen in the necessity of specification of a “sufficiently large” onset $K\gg 1$ of the de-parallelization. Such a specification is just numerically, computer-precision motivated. In contrast to the above-outlined treatment of the EPN scenarios where the dimension $N$ was fixed, the IEP-implied choice of any finite $K$ is purely pragmatic, immanently approximative and virtually arbitrary.
We now intend to show that, surprisingly enough, the apparently more or less accidental flexibility of our choice of $K$ can in fact become an important mathematical tool facilitating an EPN-resembling regularization and consistent interpretation of quantum systems near their IEP singular extreme.
First of all, there is no doubt about the necessity of transition from the less suitable unperturbed basis (composed of eigenvectors) to an “anomalous” basis resembling Eq. (4). The reason is provided by Eqs. (25) and (26): only a rectification of the underlying biorthogonal or biorthonormal basis [23] can re-establish the EPN - IEP parallels even when achieved, also in the latter case, at the expense of non-diagonality and non-Hermiticity of matrix (26).
Although the EPN singularity encountered at finite matrix dimensions $N<\infty$ is, according to paper [1], perceivably weaker than its IEP analogue, the essence of our present message will be complementary. Basically, we will emphasize that one can also reveal and make a productive use of certain partial similarities between the two concepts. In particular, we propose that the above-outlined possibility and feasibility of treating the manifestly unphysical finite-dimensional singular matrices $H^{(N)}(g^{(EPN)})$ as formally acceptable unperturbed Hamiltonians is to be transferred also to the IEP context. An anomalous “good” basis composed of the columns of transition matrix should be, mutatis mutandis, reconstructed also from any given Hamiltonian $H^{(IEP)}$.
@@ 5.2 IEP-unfolding bases
${\cal R}^{(IEP)}=\{|\psi_{0}\rangle,|\psi_{1}\rangle,\ldots,|\psi_{K-1}\rangle,|f_{K}\rangle,|f_{K+1}\rangle,|f_{K+2}\rangle,\,\ldots\}\,.$ (27)
This array is composed of the mere first $K$ eigenkets $|\psi_{j}\rangle$ complemented by the modified, associated-like ket vectors $|f_{K+k}\rangle$ with $k=0,1,2,\ldots$.
@@ 5.3 Recurrences
Our freedom of choice of the highest-component coefficients $c_{k,k}$ enables us to suppress the IEP-accompanying asymptotic parallelization of the vectors of basis in Hilbert space. The goal is achieved. For every particular IEP model we may recall recurrences (30) and replace the $Q-$projected part of the basis composed of eigenvectors by the $Q-$projected part of the basis composed, up to the first item $|f_{K}\rangle=c_{0,0}|\psi_{K}\rangle$, of non-eigenvectors. And this is precisely what has been done in Eq. (27).
@@ 6.1 Formulation of the problem
Günther with Stefani [8] stated that “what is still lacking” in the IEP setup “is a simple physical explanation scheme for the non-Rieszian behavior of the eigenfunction sets”. We agree. We are even more skeptical because we would rather say that the expected ‘simple physical explanation’ making, in particular, the popular IC oscillators (1) physical need not exist at all. Indeed, we believe that a consistent physical closed quantum system interpretation could much more easily be assigned to suitable perturbations of the “seed” IEP oscillators with uncertain interpretation (cf. also [24] in this respect).
@@
Incidentally, a similar simplification has also been detected in the realistic EPN-supporting Bose-Hubbard model of paper [12] where, in dependence on parameters, the authors had to use both  the degenerate and non-degenerate versions of perturbation theory. Thus, no abstract conceptual problems have to be expected to emerge after one returns to the generic IEP-related dynamics. Still, as long as the IEP-related problems are infinite-dimensional, the perturbed IEP spectrum cannot be deduced from any analogue of the EPN-based implicit-definition constraint (18). The study of properties of the vicinity of the IEP singularity cannot be based on a direct reference to its EPN analogue. The methods of construction have to be amended.
@@ 6.2 Structure of solutions
In the light of the $P+Q$ partitioning of matrix ${\cal J}^{(IEP)}$ in Eq. (26) the constructiton of the perturbed forms of the low-lying bound states remains standard. The $P-$projected states may be ignored just as certain decoupled observers. Only the treatment of the “asymptotic”, $Q-$projected components of the quantum system in question becomes difficult and singular, “for instance, due to spectral instabilities” [1].
@@
In our present, conceptually more ambitious analysis of the problem it makes sense to turn attention to the states with the high-lying zero-order unperturbed energies. The related unperturbed ket vectors $|\psi(0)\rangle$ will lie in the complementary (and infinite-dimensional) $Q-$projected subspace of Hilbert space. The more or less conventional construction of its perturbed descendant given by Eq. (33) will then possess several anomalous features of course.
@@
Due to the specific upper-triangular two-diagonal structure of matrix $Q{\cal J}^{(IEP)}Q$, the construction of a biorthonormalized basis would be needed. Thus, the left eigenvectors of $Q{\cal J}^{(IEP)}Q$ (i.e., in the notation of paper [20], ketkets, $|\chi_{j}\rangle\!\rangle$) will be complicated and different from their right-eigenvector biorthogonal partners $|\chi_{j}\rangle$. This means that also the conventional Rayleigh-Schrödinger elementary unperturbed projectors $|\chi_{j}\rangle\langle\!\langle\chi_{j}|$ (needed during the construction) will have a practically prohibitively complicated explicit matrix structure.
@@
Another, independent word of warning might originate from the fact that for all of the truly high energy levels $E_{K+m}(\lambda)$ with $m\gg 1$ the use of the explicit Rayleigh-Schrödinger recipe would require the derivation of formulae which would be $m-$dependent and different for the different, i.e., for the $(K+m)-$th, $m-$numbered excitations. Fortunately, the latter technical obstacle and difficulty has a comparatively elementary resolution because what is fully at our disposal is our choice of the value of $K$. We may feel free to work, exclusively, with the properly innovated Rayleigh-Schrödinger formulae deduced just at a single value of $m$, i.e., say, at $m=0$.
@@
Similarly, from the second row of Eq. (39) we would obtain the value of the second wave-function component
@@ 7 Discussion
We can conclude that in both the EPN- and IEP-related unitary-evolution scenarios the properly amended form of perturbation theory seems to be able to provide, even in its leading-order form, some explicit and useful criteria of the acceptability or unacceptability of various preselected perturbations of phenomenological interest.
@@ 7.1 Benign perturbations
Between the EPN and IEP alternatives there still exists a crucial difference. Indeed, in the typical EPN-related analysis our considerations usually start from our knowledge of the “physical” family of models $H(g)$. Then, the only truly difficult problem is to localize the EPN singularity, especially when the values of $N$ are not too small. In the case of the IEP singularities, in contrast, we only know, typically, the unperturbed Hamiltonian as sampled here by the IC operator. It is possible to conclude that precisely this makes the IEP-related models perceivably more difficult to study.
From a purely pragmatic point of view a source of certain optimism could be drawn from the leading-order perturbation-approximation criteria. Their key strength lies in the possibility of identification of the “malign” IEP perturbations which would destroy the reality of the spectrum and which would make the evolution non-unitary.
The complementary reliable identification of the “benign” perturbations is a mathematically much more difficult open problem. Incidentally, qualitatively the same conclusions have already been obtained in the simpler EPN context. For example, in the above-mentioned study [12] of a specific Bose-Hubbard model near its EPN dynamical extreme, the authors did not insist on the reality of the spectrum. They decided to treat their mathematical results as applicable and valid in a broader, not necessarily unitary open-system context.
In a narrower, closed-system setting, a deeper analysis has been performed and a resolution of the apparent EPN-related instability paradox has been described in paper [18]. We studied there the exact as well as approximate secular equations in more detail. Our ultimate conclusion was that the necessary smallness condition specifying the class of the admissible, unitarity non-violating perturbations does not involve their upper-triangular matrix part at all. In contrast, for the perturbed-EPN model in question, the lower-triangular matrix part of all of the “benign” (i.e., unitarity-compatible) perturbations has been shown to have the following element-dependent matrix form of condition of the sufficient smallness of $\lambda$,
@@ 7.2 IC oscillator as popular toy model
In order to elucidate the benchmark-model role of the IC IEP oscillator let us recall paper [5] in which Bender with Boettcher examined a rather broad family of time-independent non-Hermitian toy-model Hamiltonians (cf. Eq. (45) in Appendix A. 2 below). They felt guided by the postulate of (antilinear) symmetry of their models,
${\cal PT}H^{(BB)}=H^{(BB)}{\cal PT}\,.$ (44)
The linear operator ${\cal P}$ was treated as parity (causing the space reflection $x\to-x$) while ${\cal T}$ had to mimic the anti-linear time reversal.
The authors proposed to treat their operators $H^{(BB)}$ as “Hamiltonians whose spectra are real and positive” so that “these ${\cal PT}-$symmetric theories may be viewed as analytic continuations of conventional theories from real to complex phase space” [5]. During the subsequent wave of development of the related mathematics it has been revealed that in the language of functional analysis the ${\cal PT}-$symmetry of Eq. (44) can be re-read as pseudo-Hermiticity [14] as well as a self-adjointness in the Krein space endowed with indefinite pseudo-metric ${\cal P}$ [25, 26, 27].
@@
This was also the feature which attracted a lot of attention among physicists [5, 22, 26, 28]. Precisely because they happened to generate the purely real, discrete and non-negative (i.e., hypothetically, observable and bound-state-like) energy spectra. Still, the ultimate verdict by mathematicians [1, 8] was discouraging because they proved that the IC Hamiltonian cannot be assigned any isospectral self-adjoint avatar $\mathfrak{h}({{t}})$ or acceptable physical inner-product metric [1]. Thus, the rigorous mathematical analysis finally led to the loss of some of the most optimistic phenomenological expectations.
@@ 8 Summary
Not too surprisingly, the highly desirable proofs of the so called unbroken form of ${\cal PT}-$symmetry (in which, by definition [22], the spectrum remains real) appeared to be, in numerous applications, strongly model-dependent. There seemed to be no universal criteria guaranteeing the existence of the unbroken ${\cal PT}-$symmetry in dependence on a suitable measure of degree of the non-Hermiticity of the Hamiltonian.
@@
In our present paper we explained that and how such a decision enabled us not only to pick up a rather natural measure of the non-Hermiticity (characterized simply by the inverse distance of the variable physical parameter $g\in{\cal D}$ from its unphysical exceptional value) but also to formulate a well-defined project in which we developed and applied, consequently, some innovative and suitable perturbation-approximation techniques. In its framework we managed to show that the unbroken ${\cal PT}-$symmetry of our models can really survive inside an open parametric domain, on a point of boundary of which our measure of non-Hermiticity reaches its maximum. The latter point (which remains manifestly unphysical) has been shown to coincide either with the Kato’s [7] exceptional point of a finite order $N$ or with its hypothetical IEP analogue.
Such an approach has been found productive. Using certain slightly modified techniques of perturbation theory of linear operators with finite $N$ we found that, paradoxically, the restriction of attention to the smallest vicinity of the singularity (in which the Hamiltonians become maximally non-Hermitian) leads to a remarkable simplification of the perturbation-approximation constructions. In spite of being singular and unacceptable as observables at $g=g^{EPN}$, the special, “exceptional” non-diagonalizable operators appeared to be eligible as unperturbed Hamiltonians. In their vicinity such that $g\in{\cal D}$ their diagonalizability as well as the observability status (i.e., the standard physical status) were re-established.
In this sense, the core of our present message is that the same perturbation-regularization physical interpretation should be also attributed to the IEP models where $N=\infty$. For the purposes of illustrative example we choose the popular imaginary oscillator Hamiltonian. Such a choice has been found motivated, first of all, by its long-lasting theoretical significance which ranges from its more or less purely formal role in mathematics and functional analysis [55] up to a deeper phenomenological significance in quantum statistics (where the imaginary $\phi^{3}$ interaction mimics the Yang-Lee edge singularity [5, 21]) and up to its important theoretical role of a benchmark model in the conformal quantum field theory [56] as well as in the less well known but still popular Reggeon field theory [57].
In all of these contexts our present results imply that the IEP property of the IC-like models means unphysicality. Only a suitable perturbation can reinstall the (possibly, “hidden” alias  “quasi-”) unitarity and physicality. Thus, the practical realizations of the standard quantum-mechanical IC model remain, in a way and for the reasons as outlined in [1], elusive, in the unitary-theory context at least [35]. At the same time one might still expect that some of its realizations could emerge off the realm of quantum mechanics, i.e., say, in optics [29].
Constructively we managed to clarify also some of the consequences of our present perturbative-regularization recipe. The emergence of qualitative as well as quantitative EPN - IEP parallels helped us to complement and understand better the twelve years old disproof of the internal mathematical consistency of the IC IEP quantum oscillator [1]. Such a clarification can be perceived as being of a fundamental importance in quantum theory. Indeed, potentially, most of our observations might immediately be extended also to many other currently popular but IEP-singular non-Hermitian quantum models.
@@
Fortunately, an ultimate resolution of the conflict has been found in a rediscovery and return to a half-forgotten but still fully relevant older review paper by Scholtz et al [10]. In it, most of the objections by mathematicians were circumvented by an ad hoc  technical assumption that one is only allowed to consider the non-Hermitian Hamiltonians (as well as any other candidates for observables) which are, as operators in Hilbert space, bounded: cf. also several mathematically oriented reviews in [26] in this respect. Thus, one may call such a mathematically consistent version of the theory quasi-Hermitian quantum mechanics.